OpenStudy (anonymous):

determine the equation of the line tangent to the graph of y=15e^(6x^2)-6x at x=2 in the form y=mx+b i got m=2160e^144 b=-4305e^144-12 but those are wrong

OpenStudy (anonymous):

find the first derivative and then put x=2 you will get the slop at x=2 \[y \prime = 15 e ^{6x ^{2}}\left( 12x \right) - 6\] now put x=2

OpenStudy (anonymous):

15e^(144)(24)-6?

OpenStudy (anonymous):

or would it be 15e^(24)(24)-6

OpenStudy (anonymous):

first yu have x = 2, now sub 2 where ever x is in the original equation and that gives you the y value.

OpenStudy (anonymous):

second, take the first derivative of the original function then sub x = 2, which is gonna give you the m :) and then yu can plug in y-y1=m(x-x1) and that's it :)

OpenStudy (anonymous):

so is my answer i found above that i put in the question completely wrong?

OpenStudy (anonymous):

i'll check

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait actually i think y should equal 15e^(144)-12

OpenStudy (anonymous):

no, it wud be y = 15e^(24) - 12

OpenStudy (anonymous):

oo i multiplied the order wrong

OpenStudy (anonymous):

what m=?

OpenStudy (anonymous):

did u get this for the derivative?

OpenStudy (anonymous):